# Integral canonical models of exceptional Shimura varieties

### v1

@@ Abstract
We prove that Shimura varieties admit integral canonical models for sufficiently large primes. In the case of abelian-type Shimura varieties, this recovers work of Kisin-Kottwitz for sufficiently large primes. We also prove the existence of integral canonical models for images of period maps corresponding to geometric families. We deduce several consequences from this, including an unramified rigid analogue of Borel’s extension theorem, a version of Tate semisimplicity, CM lifting theorems, and a version of Tate’s isogeny theorem for ordinary points.
@@ 1 Introduction
Langlands suggested more generally that integral canonical models should exist whenever the compact open $K_{p}$ at $p$ is hyperspecial. In the case of PEL Shimura varieties (i.e. Shimura varieties parameterizing polarized abelian varieties with extra endomorphisms), Kottwitz in [Ko] constructs integral models when the level structure at $p$ is hyperspecial. Kisin generalizes Kottwitz’ work to the setting of Shimura varieties of abelian type ([Kisinintegral]). In [PappasRapoport], Pappas and Rapoport lay out a framework for integral models for all Shimura varieties, and amongst various other things, they show that the integral models constructed by Kisin satisfy their framework. In [Lovering], Lovering constructs integral canonical models for automorphic data associated to Shimura varieties of abelian type. The extension property proven in most of these works ultimately relies on the work of Chai and Faltings, which uses the moduli intepretation of $\mathcal{A}_{g}$, and therefore does not carry over to exceptional Shimura varieties.
@@ 1.2.1 Applications
We deduce several applications for mod $v$ (and $v$-adic) Shimura varieties.
Firstly, we prove the following unramified $p$-adic analogue of Borel’s theorem, which we will state in the proper case for ease of exposition (see Section 6 for a more general statement)
Theorem 1.4.
Let $S_{K}(G,X)$ be an exceptional proper Shimura variety. Let $\mathsf{D}$ denote the closed unit rigid-analytic disc, and let $\mathsf{D}^{\times}=\mathsf{D}\setminus\{0\}$. Let $f:\mathsf{D}^{\times}\rightarrow S_{K}(G,X)^{{\mathrm{an}}}$ denote a rigid-analytic map defined over $W(\overline{\mathbb{F}}_{p})[1/p]$. Then, $f$ extends to a map $\mathsf{D}\rightarrow S_{K}(G,X)^{{\mathrm{an}}}$.
In fact we also prove a version of Theorem 1.4 in the context of integral models of images of period maps (see Theorem 1.9).
We also prove results pertaining to the mod $v$ arithmetic and geometry of $\mathscr{S}_{K}(G,X)$. The main theorem implies that Hecke correspondences away from $p$ extend to $\mathscr{S}_{K}(G,X)/\mathcal{O}_{E_{v}}$. We use these Hecke correspondences to prove the following theorem:
@@
1.3 Comparison with the extension property of Milne-Moonen
@@ 1.4 Integral canonical models for images of geometric period maps
Now let $P$ be an irreducible normal variety defined over a number field $E$, $f:Z\to P$ a smooth projective $F$-morphism, and for any fixed $k$ let ${}_{et}\mathbb{W}_{\ell}:=R^{k}f_{*}\mathbb{Z}_{\ell}$ be the arithmetic local system—or more generally a subquotient obtained from maps between such local systems induced by $P$-morphisms of families [^1] We lay out the precise conditions for a variation of Hodge structures to “come from geometry” in the sense we need it in Section 4.. We let $Y$ be the Stein factorization of the period map of the variation with underlying geometric local system $\mathbb{V}_{\mathbb{Z}}=R^{k}f^{\mathrm{an}}_{*}\mathbb{Z}$ for some embedding $E\subset\mathbb{C}$. Then $Y$ is defined over $E$, does not depend on the embedding, and the local system ${}_{et}\mathbb{W}_{\ell}$ descends to $Y$.
@@
Let $E$ be a number field. Let $Y/F$ be the Stein factorization of the period map of a geometric variation of Hodge structures which we assume to be inertia-free (and therefore a quasiprojective variety). Let ${}_{et}\mathbb{V}_{\ell}$ be the descent of the natural $\ell$-adic arithmetic local system to $Y$.
@@
      For any irreducible smooth $\mathcal{X}/\mathcal{O}_{v}$ with generic point $\eta_{\mathcal{X}}$, any morphism $\phi_{0}:\eta_{\mathcal{X}}\to Y_{E_{v}}$ extends to $\phi:\mathcal{X}\to\mathcal{Y}_{v}$ if and only if $\phi_{0}^{*}(_{et}\mathbb{V}_{\ell})$ extends to $\mathcal{X}$ for some $\ell\neq p$.
    3. (c)
      There exists a global model $\mathcal{Y}$ of $Y$ which restricts to $\mathcal{Y}_{v}$ for all places $v$ of $\mathcal{O}_{E,\Sigma_{Y}}$.
@@
    For any $\mathcal{O}_{v}$-model $\mathcal{Z}_{v}$ of $Y$ that admits a uniform log smooth resolution (see Definition 2.12) and to which ${}_{et}\mathbb{V}_{\ell}$ extends for some $\ell\neq p$, the identity map over the generic fiber extends to a $\mathcal{O}_{v}$-morphism $\mathcal{Z}_{v}\to\mathcal{Y}_{v}$.
@@
    Let $Y,Y^{\prime}/E$ be the Stein factorizations of two possibly different period maps arising as above, and for a finite place $v$ of $\mathcal{O}_{E,\Sigma_{Y}\cup\Sigma_{Y^{\prime}}}$ let $\mathcal{Y}_{v},\mathcal{Y}^{\prime}_{v}$ be models as in (1). Then any morphism $\phi_{0}:\eta_{\mathcal{Y}}\to\mathcal{Y}^{\prime}_{v}$ extends to $\phi:\mathcal{Y}_{v}\to\mathcal{Y}^{\prime}_{v}$ if and only if $\phi_{0}^{*}(_{et}\mathbb{V}_{\ell})$ extends to $\mathcal{Y}_{\mathcal{O}_{v}}$ for some $\ell\neq p$. In particular, the local model $\mathcal{Y}_{v}$ of $Y$ is canonically determined.
In general, the image of a period map has no reason to be smooth (already over $\mathbb{C}$), so it is important to allow for singularities. Even starting with a smooth moduli space $S$ parametrizing varieties with an infinitesimal Torelli theorem so that the period map $S^{\mathrm{an}}\to[\mathbf{G}(\mathbb{Z})\backslash D]$ is immersive, it will often not be proper; the “Hodge-theoretic” compactification is essentially the closure of the image of $S^{\mathrm{an}}\to\mathbf{G}(\mathbb{Z})\backslash D$ (up to a finite map) and will therefore usually be singular anyway. This is already true for the image of the Torelli map for the moduli space of curves (see 2.11).
1.5 A purity property for Fontaine–Laffaille modules
The extension property in the settings of Theorem 1.3 and Theorem 1.9 boils down to a purity property of a $p$-torsion Fontaine–Laffaille module pulled back from a variety on which it has immersive Kodaira–Spencer map. It is motivated by classical Hodge theory and the arguments of [dejongoort, moretbailly, faltingschai]. Briefly, let $F$ be a mixed characteristic local field with ring of integers $\mathcal{O}$ and residue field $k$, $\mathcal{X}/\mathcal{O}$ a smooth scheme, and $f_{F}:X\to Y$ a morphism over the generic point to a model of an image of a period map $\mathcal{Y}$. There are three steps to proving the extension of $f_{F}$ to a morphism $f:\mathcal{X}\to\mathcal{Y}$:
@@ 1.5 A purity property for Fontaine–Laffaille modules
    The main step is then to handle the case $\dim\mathcal{X}=2$, and here we must use that $\mathcal{Y}$ supports a crystalline local system whose associated Fontaine–Laffaille module has immersive Kodaira–Spencer map mod $p$. In this case, $\mathcal{X}/\mathcal{O}$ is a family of smooth curves, and can be extended on a blow-up $\mathcal{X}^{\prime}$ at points in the special fiber. It is proven in [OV] that the strong divisibility property of $p$-torsion Fontaine–Laffaille modules implies a stability property: they cannot contain positive degree subsheaves. Fontaine–Laffaille modules cannot be pulled back along maps from nodal curves if the lift to $W_{2}$ smooths out some nodes, since there are no longer local lifts of Frobenius and therefore the inverse Cartier transform no longer makes sense. We show however it exists in a weaker sense up to twisting by the nodes, and this is enough to rule out nodal curves in $\mathcal{Y}$ that lift to $W_{2}$ and smooth in a controlled way. In particular it follows that the exceptional curves must be contracted—otherwise, the positive degree tangent bundle of a rational curve would contradict the stability property we prove. For a precise statement, see Proposition 3.9.
@@ 1.6 Comparison with past work
Regarding the applications to the mod $p$ and $p$-adic geometry of Shimura varieties, Kisin ([Kisinmodppoints]) proves that every $\overline{\operatorname{\mathbb{F}}}_{p}$-point of abelian Shimura varieties admit CM lifts up to isogeny. Note that the up-to-isogeny requirement is crucial as work of Oort [oortJAG] shows (for a more general statement, see also [KLSS]). Noot in [Noot] proves the existence of CM lifts (on the nose) for points of Hodge-type Shimura varieties that intersect the ordinary locus of $\mathcal{A}_{g}$. Noot’s theorem was generalized to $\mu$-ordinary points by [MoonenSerreTate] in the PEL-type case to the and [ShankarZhou] in the Hodge type case. In [padicBorel], the authors prove the $p$-adic extension and algebraicity theorems for maps to Shimura varieties (and Rapoport-Zink spaces) of abelian type defined over discretely valued $p$-adic fields and where the source of the map can have any dimension.
@@ 1.7 Acknowledgements
B. B. benefited from many conversations with Yohan Brunebarbe throughout, and first learned of the work of Ogus–Vologodsky from Raju Krishnamoorthy. We are also grateful to Michael Groechenig, Kiran Kedlaya, Mark Kisin, Arthur Ogus, Abhishek Oswal, and Vadim Vologodsky for helpful conversations. B. B. was partially supported by NSF DMS-2131688. A. S. was partially supported by the NSF grant DMS-2100436, and would like thank MSRI for their hospitality during the spring of 2023.
@@ 1.8 Outline
In Section 2 we formulate the definition of an integral canonical model in terms of the extension property, and show that extension in codimension 1 can be controlled in many cases (including those of interest) by a local system. We also discuss the descent and extension of arithmetic local systems to integral models of period images, and the relation of the extension property to purity results in classical Hodge theory. In Section 3 we prove some weak semistability results for the pull-backs of Fontaine–Laffaille modules to nodal curves. In Section 4 we formulate precise conditions under which the presence of a Fontaine–Laffaille module on a local model implies the extension property, and in Section 5 we apply it to the case of Shimura varieties and period images to prove the existence of integral canonical models as in Theorem 1.3 and Theorem 1.9. In Section 6 we use the same semistability results to prove Theorem 1.4. In Section 7 we prove Hecke correspondences extend and use this to prove Theorem 1.5. In Sections 8 and 9, we analyze the complete local rings of the integral canonical models of Shimura varieties at closed points and prove Theorems 1.6, 1.7, and 1.8.
@@ 1.9 Notation
With the exception of Section 3, we use unadorned Roman letters $S,Y$ to denote algebraic varieties over a field $F$ (typically a number field or a characteristic 0 local field), script letters $\mathcal{S},\mathcal{Y}$ to denote their integral models over a ring of integers $\mathcal{O}_{F}$ in $F$, and a subscript (either $\mathcal{S}_{v},\mathcal{Y}_{v}$ or $\mathcal{S}_{0},\mathcal{Y}_{0}$) to denote the special fiber—the base-change to the residue field $k$ of $\mathcal{O}_{F}$—at a finite place $v$ (or 00 in the local situation, when there is no ambiguity). In section Section 3, we denote by unadorned Roman letters $S,Y$ varieties over a perfect characteristic $p$ field $k$, and by $\tilde{S},\tilde{Y}$ a lift to $W_{2}(k)$.
@@ 2 The extension property and natural boundary
In this section we single out a class of integral models for which extension on the level of fundamental groups is the only obstruction to extending morphisms defined over the generic point. In particular, such models will be uniquely determined when they exist. The key property these models will have is the extension of morphisms from smooth schemes, just as in the work of Milne and Moonen. In the Shimura case (or whenever the model is smooth) this property obviously uniquely determines the model; in the general case it will be important to allow for singularities, as even for smooth moduli spaces, the extension property will often force a partial compactification which is not smooth—for instance, see 2.11. We also discuss how the extendability on the level of fundamental groups can be detected by the extendability of local systems.
@@ 2.1 The extension property
Let $\mathcal{O}$ be a complete DVR with fraction field $F$ and residue field $k$.
@@ 2.2 Natural boundary and extension in codimension 1
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$.
@@
    Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme. A resolution of $\mathcal{Y}/\mathcal{O}$ is a smooth $\mathcal{O}$-scheme $\mathcal{S}/\mathcal{O}$ with a proper dominant map $\pi:\mathcal{S}\to\mathcal{Y}$ such that $\mathcal{O}_{\mathcal{Y}}\xrightarrow{\cong}\pi_{*}\mathcal{O}_{\mathcal{S}}$ via the natural map.
@@
  4. 4.
    Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme, $\pi:\mathcal{S}\to\mathcal{Y}$ a resolution, and $\bar{\mathcal{S}}/\mathcal{O}$ a log smooth compactification of $\mathcal{S}$. We say that the boundary of $\bar{\mathcal{S}}$ is natural (with respect to $\mathcal{Y}$) if for every generic point $\eta$ of a stratum of $\mathcal{D}$, the geometric inertia of $\eta$ injects in $\pi_{1}^{{\mathrm{et}}}(\mathcal{Y},\bar{y})$ for a chosen geometric basepoint $\bar{y}$ of $\mathcal{Y}$.
@@
Let $\mathcal{O}$ be a characteristic 0 complete DVR with fraction field $F$. Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme, $\pi:\mathcal{S}\to\mathcal{Y}$ a resolution, and $\bar{\mathcal{S}}/\mathcal{O}$ a log smooth compactification of $\mathcal{S}$ with natural boundary. Let $\mathcal{X}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme and $f_{F}:X\to Y$ an $\mathcal{O}$-admissible morphism over the generic point. Then there is a finite type flat $\mathcal{X}^{\prime}/\mathcal{O}$, a proper dominant generically finite morphism $\mathcal{X}^{\prime}\to\mathcal{X}$, and a morphism $\mathcal{X}^{\prime}\to\mathcal{S}$ such that the following diagram on the generic fibers commutes:
@@
The map $f_{F}:X\to Y$ lifts to $S$ over a dense open subset of $X$; let $\mathcal{W}\subset\mathcal{X}\times_{\mathcal{O}}\bar{\mathcal{S}}$ be the closure of the graph. We then have that $\mathcal{W}\to\mathcal{X}$ is proper dominant with a morphism $\mathcal{W}\to\bar{\mathcal{S}}$, and taking an appropriate complete intersection of ample hypersurfaces, we may find $\mathcal{X}^{\prime}\subset\mathcal{W}$ which is proper dominant generically finite over $\mathcal{X}$. If $X$ meets the locus where $\pi$ is an isomorphism, then we take $\mathcal{X}^{\prime}=\mathcal{W}$ and the morphism $\mathcal{X}^{\prime}\to\mathcal{X}$ is a generic isomorphism. Either way, if the boundary of $\bar{\mathcal{S}}$ intersects $\mathcal{X}^{\prime}$ nontrivially, then there is a finite extension $\mathcal{O}^{\prime}$ of $\mathcal{O}$ an $\mathcal{O}^{\prime}$-point of $\mathcal{X}^{\prime}/\mathcal{O}$ whose closed point maps to the boundary. Because the boundary is natural, by the argument of [PST, Thm 4.4] this contradicts the admissibility assumption.

### v2

@@ Abstract
We prove that Shimura varieties admit integral canonical models for sufficiently large primes. In the case of abelian-type Shimura varieties, this recovers work of Kisin-Kottwitz for sufficiently large primes. We also prove the existence of integral canonical models for images of period maps corresponding to geometric families. We deduce several consequences from this, including a version of Tate semisimplicity, CM lifting theorems, and a version of Tate’s isogeny theorem for ordinary points.
@@ 1 Introduction
Langlands suggested more generally that integral canonical models should exist whenever the compact open $K_{p}$ at $p$ is hyperspecial. In the case of PEL Shimura varieties (i.e. Shimura varieties parameterizing polarized abelian varieties with extra endomorphisms), Kottwitz in [Ko] constructs integral models when the level structure at $p$ is hyperspecial. Kisin generalizes Kottwitz’ work to the setting of Shimura varieties of abelian type ([Kisinintegral]). In [imai2023prismatic] the authors prove a prismatic characterization of integral models of abelian type. In [PappasRapoport], Pappas and Rapoport lay out a framework for integral models for all Shimura varieties, and amongst various other things, they show that the integral models constructed by Kisin satisfy their framework. In [Lovering], Lovering constructs integral canonical models for automorphic data associated to Shimura varieties of abelian type. The extension property proven in most of these works ultimately relies on the work of Chai and Faltings, which uses the moduli intepretation of $\mathcal{A}_{g}$, and therefore does not carry over to exceptional Shimura varieties.
@@ 1.2.1 Applications
We deduce several applications for mod $v$ (and $v$-adic) Shimura varieties. The main theorem implies that Hecke correspondences away from $p$ extend to $\mathscr{S}_{K}(G,X)/\mathcal{O}_{E_{v}}$. We use these Hecke correspondences to prove the following theorem:
@@
1.3 Comparison with the extension property of Milne–Moonen
@@ 1.4 Integral canonical models for images of geometric period maps
Now let $P$ be an irreducible normal variety defined over a number field $E$, $f:Z\to P$ a smooth projective $E$-morphism, and for any fixed $k$ let ${}_{et}\mathbb{W}_{\ell}:=R^{k}f_{*}\mathbb{Z}_{\ell}$ be the arithmetic local system—or more generally a subquotient obtained from maps between such local systems induced by $P$-morphisms of families [^1] We lay out the precise conditions for a variation of Hodge structures to “come from geometry” in the sense we need it in Section 4.. We let $Y$ be the Stein factorization of the period map of the variation with underlying geometric local system $\mathbb{V}_{\mathbb{Z}}=R^{k}f^{\mathrm{an}}_{*}\mathbb{Z}$ for some embedding $E\subset\mathbb{C}$. Then $Y$ is defined over $E$, does not depend on the embedding, and the local system ${}_{et}\mathbb{W}_{\ell}$ descends to $Y$.
@@
Let $E$ be a number field. Let $Y/E$ be the Stein factorization of the period map of a geometric variation of Hodge structures which we assume to be inertia-free (and therefore a quasiprojective variety). Let ${}_{et}\mathbb{V}_{\ell}$ be the descent of the natural $\ell$-adic arithmetic local system to $Y$.
@@
      For any irreducible smooth $\mathcal{X}/\mathcal{O}_{v}$ with generic point $\eta_{\mathcal{X}}$, any morphism $g_{0}:\eta_{\mathcal{X}}\to Y_{E_{v}}$ extends to $g:\mathcal{X}\to\mathcal{Y}_{v}$ if and only if $g_{0}^{*}(_{et}\mathbb{V}_{\ell})$ extends to $\mathcal{X}$ for some $\ell\neq p$.
@@
    The models in part 1 are functorial in the following sense. Let $Y,Y^{\prime}/E$ be the Stein factorizations of two possibly different period maps arising as above, and for a finite place $v$ of $\mathcal{O}_{E,\Sigma_{Y}\cup\Sigma_{Y^{\prime}}}$ let $\mathcal{Y}_{v},\mathcal{Y}^{\prime}_{v}$ be models as in part 1. Then any morphism $g_{0}:\eta_{\mathcal{Y}}\to\mathcal{Y}^{\prime}_{v}$ extends to $g:\mathcal{Y}_{v}\to\mathcal{Y}^{\prime}_{v}$ if and only if $g_{0}^{*}(_{et}\mathbb{V}_{\ell})$ extends to $\mathcal{Y}_{v}$ for some $\ell\neq p$. In particular, the local model $\mathcal{Y}_{v}$ of $Y$ is canonically determined.
@@
    The models in part 1 are minimal in the following sense. For any $\mathcal{O}_{v}$-model $\mathcal{Z}_{v}$ of $Y$ that admits a uniform log smooth resolution (see Definition 2.13) and to which ${}_{et}\mathbb{V}_{\ell}$ extends for some $\ell\neq p$, the identity map over the generic fiber extends to an $\mathcal{O}_{v}$-morphism $\mathcal{Z}_{v}\to\mathcal{Y}_{v}$.
  4. 4.
    Finally, there exists a global model $\mathcal{Y}$ over $\mathcal{O}_{E,\Sigma_{Y}}$ of $Y$ which restricts to the model $\mathcal{Y}_{v}$ of part 1 for all places $v$ of $\mathcal{O}_{E,\Sigma_{Y}}$.
In general, the image of a period map has no reason to be smooth (already over $\mathbb{C}$), so it is important to allow for singularities. Even starting with a smooth moduli space $S$ parametrizing varieties with an infinitesimal Torelli theorem so that the period map $S^{\mathrm{an}}\to[\mathbf{G}(\mathbb{Z})\backslash D]$ is immersive, it will often not be proper; the “Hodge-theoretic” compactification is essentially the closure of the image of $S^{\mathrm{an}}\to[\mathbf{G}(\mathbb{Z})\backslash D]$ (up to a finite map) and will therefore usually be singular anyway. This is already true for the image of the Torelli map for the moduli space of curves (see 2.12).
1.5 An outline of the proof of the Extension Property
The proof of the extension property in Theorem 1.3 is motivated by classical Hodge theory and the arguments of [dejongoort, moretbailly, faltingschai]. Briefly, let $F/\mathbb{Q}_{p}$ be a finite unramified extension ring of integers $\mathcal{O}$, $\mathcal{X}/\mathcal{O}$ a smooth scheme, and $f_{F}:X\to Y$ a morphism over the generic point to a model of a Stein factorization of a period map $\mathcal{Y}$. There are three steps to proving the extension of $f_{F}$ to a morphism $f:\mathcal{X}\to\mathcal{Y}$:
@@ 1.5 An outline of the proof of the Extension Property
    The main step is then to handle the case $\dim\mathcal{X}=2$, so when $\mathcal{X}/\mathcal{O}$ is a family of smooth curves, and here we must use that $\mathcal{Y}$ supports [^2] This only makes sense after pulling back to a resolution which is smooth over $\mathcal{O}$ a crystalline local system whose associated Fontaine–Laffaille module has ample Griffiths bundle. On the one hand, by a key result of Hokaj [Hokaj] and Guo–Yang [guo], the Fontaine–Laffaille module extends to $\mathcal{X}/\mathcal{O}$; on the other hand, the map to $\mathcal{Y}$ can be extended on a blow-up $\mathcal{X}^{\prime}$ at points in the special fiber. The pull-backs of the two Fontaine–Laffaille modules must agree, and the ampleness of the Griffiths bundle implies every exceptional curve is contracted in $\mathcal{Y}$.
@@ 1.6 Comparison with past work
Regarding the applications to the mod $p$ and $p$-adic geometry of Shimura varieties, Kisin ([Kisinmodppoints]) proves that every $\overline{\operatorname{\mathbb{F}}}_{p}$-point of abelian Shimura varieties admit CM lifts up to isogeny. Note that the up-to-isogeny requirement is crucial as work of Oort [oortJAG] shows (for a more general statement, see also [KLSS]). Noot in [Noot] proves the existence of CM lifts (on the nose) for points of Hodge-type Shimura varieties that intersect the ordinary locus of $\mathcal{A}_{g}$. Noot’s theorem was generalized to $\mu$-ordinary points by [MoonenSerreTate] in the PEL-type case to the and [ShankarZhou] in the Hodge type case.
@@ 1.7 Acknowledgements
B. B. benefited from many conversations with Yohan Brunebarbe throughout, and first learned of the work of Ogus–Vologodsky from Raju Krishnamoorthy. We are also grateful to Michael Groechenig, Kiran Kedlaya, Mark Kisin, Arthur Ogus, Abhishek Oswal, and Vadim Vologodsky for helpful conversations. We are particularly indebted to Alex Youcis for pointing out the result of Hokaj [Hokaj], which enabled us to correct an error in the original version. B. B. was partially supported by NSF DMS-2131688. A. S. was partially supported by the NSF grant DMS-2100436 and DMS-2338942, and would like thank MSRI for their hospitality during the spring of 2023.
@@ 1.8 Outline
In Section 2 we formulate the definition of an integral canonical model in terms of the extension property, and show that extension in codimension 1 can be controlled in many cases (including those of interest) by a local system. We also discuss the descent and extension of arithmetic local systems to integral models of period images, and the relation of the extension property to purity results in classical Hodge theory. In Section 3 we prove the extension in codimension 2. In Section 4 we formulate precise conditions under which the presence of a Fontaine–Laffaille module on a local model implies the extension property, and in Section 5 we apply it to the case of Shimura varieties and period images to prove the existence of integral canonical models as in Theorem 1.3 and Theorem 1.8. In Section 6 we prove Hecke correspondences extend and use this to prove Theorem 1.4. In Sections 7 and 8, we analyze the complete local rings of the integral canonical models of Shimura varieties at closed points and prove Theorems 1.5, 1.6, and 1.7.
@@ 1.9 Notation
Typically $\mathcal{O}$ will be the ring of integers in either a number field $E$ or an unramified finite extension $F/\mathbb{Q}_{p}$. We use script letters $\mathcal{S},\mathcal{X},\mathcal{Y}$ to denote either finite type $\operatorname{Spec}\mathcal{O}$-schemes or finite type formal $\operatorname{Spf}\mathcal{O}$-scheme (in the latter case). We denote by unadorned Roman letters $S,X,Y$ the generic fibers over $E$ or $F$ (in the non-formal case) and $\mathcal{S}_{0},\mathcal{X}_{0},\mathcal{Y}_{0}$ will denote the special fibers (when $\mathcal{O}$ is local). For a finite place $v$ of $\mathcal{O}_{E}$ we denote by $\mathcal{O}_{v}=\mathcal{O}_{E_{v}}$ the completion.
@@ 2 The extension property and natural boundary
In this section we single out a class of integral models for which extension on the level of fundamental groups is the only obstruction to extending morphisms defined over the generic point. In particular, such models will be uniquely determined when they exist. The key property these models will have is the extension of morphisms from smooth schemes, just as in the work of Milne and Moonen. In the Shimura case (or whenever the model is smooth) this property obviously uniquely determines the model; in the general case it will be important to allow for singularities, as even for smooth moduli spaces, the extension property will often force a partial compactification which is not smooth—for instance, see 2.12. We also discuss how the extendability on the level of fundamental groups can be detected by the extendability of local systems.
Throughout this section, we let $\mathcal{O}$ be a Dedekind domain with fraction field $F$ which is either a finite extension of $\mathbb{Q}_{p}$ or a global number field. In the local situation, we denote by $k$ the residue field.
@@ 2.2 Natural boundary and extension in codimension 1
    Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme. A resolution of $\mathcal{Y}/\mathcal{O}$ is a smooth $\mathcal{O}$-scheme $\mathcal{S}/\mathcal{O}$ with a proper dominant birational map $\pi:\mathcal{S}\to\mathcal{Y}$ such that $\mathcal{O}_{\mathcal{Y}}\xrightarrow{\cong}\pi_{*}\mathcal{O}_{\mathcal{S}}$ via the natural map.
@@
In the cases that will be of interest to us, the boundary of $\bar{\mathcal{S}}$ will be detected by the fundamental group in the sense made precise by the following definition. For a log smooth scheme $(\bar{S},D)$ over $\mathbb{C}$ with $S=\bar{S}\setminus D$ and whose boundary $D$ has irreducible components $\{D_{i}\}_{i\in I}$, and for any point $x\in D$ together with a small arc $\hat{x}:[0,1)\to\bar{S}$ mapping 00 to $x$ and $(0,1)$ to $S$, the local fundamental group $\pi_{1}^{\mathrm{loc}}(\bar{S},D,\hat{x}):=\lim_{x\in U\subset\bar{S}}\pi_{1}(U\setminus D,\hat{x})$ is naturally identified with $\bigoplus_{D_{i}\ni x}\mathbb{Z}$, where the sum is over irreducible component $D_{i}$ of $D$ containing $x$ and the basepoint is taken to be any point on $\hat{x}$ sufficiently close to $x$.
Definition 2.3.
Let $(\bar{S},D)$ be a proper log smooth complex variety with $S=\bar{S}\setminus D$ and $\pi:S\to Y$ a resolution of a normal complex variety $Y$. We say the boundary of $\bar{S}$ is $\pi$-natural if for every morphism $f:(\bar{C},D_{C})\to(\bar{S},D)$ from a log smooth proper curve $(\bar{C},D_{C})$ with $D_{C}=(f^{*}D)^{\mathrm{red}}$ and for any $c\in D_{C}$, we have that the composition
$\pi_{1}^{\mathrm{loc}}(C,D_{C},\hat{c})\to\pi_{1}^{\mathrm{loc}}(\bar{S},D,f(c))\to\pi_{1}(S,f(\hat{c}))\to\pi_{1}(Y,\pi\circ f(\hat{c}))\to\pi_{1}^{\mathrm{et}}(Y,\pi\circ f(\hat{c}))$
is injective. Equivalently, we require that for any $f,c$ as above and any $n$, there is a Galois finite étale cover $Y^{\prime}\to Y$ whose base-change to $C$ ramifies to order $\geq n$ at $c$.
Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme, $\pi:\mathcal{S}\to\mathcal{Y}$ a resolution, and $\bar{\mathcal{S}}/\mathcal{O}$ a log smooth compactification of $\mathcal{S}$. We say that the boundary of $\bar{\mathcal{S}}$ is $\pi$-natural if this is true for the base-change of $\pi$ and $\bar{\mathcal{S}}$ to $\mathbb{C}$ for some (hence any, by the above remark) embedding of $K$ into $\mathbb{C}$. If $\pi$ is the identity we just say $\bar{\mathcal{S}}$ has natural boundary.
@@
Let $F/\mathbb{Q}_{p}$ be an unramified extension with ring of integers $\mathcal{O}$. Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme, $\pi:\mathcal{S}\to\mathcal{Y}$ a resolution, and $\bar{\mathcal{S}}/\mathcal{O}$ a log smooth compactification of $\mathcal{S}$ with $\pi$-natural boundary. Let $\mathcal{X}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme and $f_{F}:X\to Y$ an $\mathcal{O}$-admissible morphism over the generic point. Then there is a finite type flat $\mathcal{X}^{\prime}/\mathcal{O}$, a proper dominant generically finite morphism $\mathcal{X}^{\prime}\to\mathcal{X}$, and a morphism $\mathcal{X}^{\prime}\to\mathcal{S}$ such that the following diagram on the generic fibers commutes:
@@
Let $\mathcal{W}\subset\mathcal{X}\times_{\mathcal{O}}\bar{\mathcal{S}}$ be the closure of the fiber product $X\times_{Y}S\subset\mathcal{X}\times_{\mathcal{O}}\bar{\mathcal{S}}$. We then have that the projection $\mathcal{W}\to\mathcal{X}$ is proper dominant with a morphism $\mathcal{W}\to\bar{\mathcal{S}}$, and taking an appropriate complete intersection of ample hypersurfaces, we may find $\mathcal{X}^{\prime}\subset\mathcal{W}$ which is proper dominant generically finite over $\mathcal{X}$. If $X$ meets the locus where $\pi$ is an isomorphism, then we take $\mathcal{X}^{\prime}=\mathcal{W}$ and the morphism $\mathcal{X}^{\prime}\to\mathcal{X}$ is a generic isomorphism. Either way, if the boundary of $\bar{\mathcal{S}}$ intersects $\mathcal{X}^{\prime}$ nontrivially, then there is a finite extension $\mathcal{O}^{\prime}$ of $\mathcal{O}$ and an $\mathcal{O}^{\prime}$-point of $\mathcal{X}^{\prime}/\mathcal{O}$ whose closed point maps to the boundary. Because the boundary is natural, by the argument of [PST, Thm 4.4] this contradicts the admissibility assumption.